OpenStudy (anonymous):

give the range of x -2y=0

OpenStudy (therealmeeeee):

x - 1 > 0 x > 1 y > 2 x-int: ln (x - 1) + 2 = 0 ln (x - 1) = -2 x - 1 = e^(-2) x = 1 + e^(-2) no y-int.

OpenStudy (therealmeeeee):

do this help

OpenStudy (therealmeeeee):

omg dude do this help